# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

### v2

@@ 3.3.3. Operation Phase
The operation phase is where the MCP server actively processes requests, executes tool invocations, and facilitates seamless interaction between AI applications and external resources. Tool execution allows the MCP server to invoke the appropriate tools based on the AI application’s requests, ensuring that the selected tools perform their intended operations. Slash command handling enables the server to interpret and execute multiple commands, including those issued through user interfaces or AI agents, while managing potential command overlaps to prevent conflicts. Sandbox mechanism enforcement ensures that the execution environment is isolated and secure, preventing unauthorized access and mitigating potential risks. Throughout the operation phase, the MCP server maintains a stable and controlled environment, enabling reliable and secure task execution.
3.3.4. Update Phase
The update phase ensures that the MCP server remains secure, up-to-date, and capable of adapting to evolving requirements. This phase includes three key tasks. Authorization management verifies that post-update access permissions remain valid, preventing unauthorized use of server resources after updates. Version control maintains consistency between different server versions, ensuring that new updates do not introduce vulnerabilities or conflicts. Old version management deactivates or removes outdated versions to prevent attackers from exploiting known vulnerabilities in previous versions.
Understanding the MCP server lifecycle is essential for identifying potential vulnerabilities and designing effective security measures. Each phase introduces distinct challenges that must be carefully addressed to maintain the security, efficiency, and adaptability of the MCP server in dynamic AI environments.
@@ 4.1. Ecosystem Overview
Table 1 demonstrates how MCP has gained significant traction across diverse sectors, signaling its growing importance in enabling seamless AI-to-tool interactions. Notably, leading AI companies such as Anthropic (Anthropic, 2024a) and OpenAI (OpenAI, 2025) have integrated MCP to enhance agent capabilities and improve multi-step task execution. This adoption by industry pioneers has set a precedent, encouraging other major players to follow suit. Chinese tech giants like Baidu (Maps, 2025) have also incorporated MCP into their ecosystems, highlighting the protocol’s potential to standardize AI workflows across global markets. Developer tools and IDEs, including Replit (Replit, 2025), Microsoft Copilot Studio (Studio, 2025), JetBrains (JetBrains, 2025), and TheiaIDE (TheiaAI/TheiaIDE, 2025), leverage MCP to facilitate agentic workflows and streamline cross-platform operations. This trend indicates a shift toward embedding MCP in developer environments to enhance productivity and reduce manual integration efforts. Furthermore, cloud platforms like Cloudflare (Cloudflare, 2025) and financial service providers such as Block (Square) ((2025), Square) and Stripe (Stripe, 2025) are exploring MCP to improve security, scalability, and governance in multi-tenant environments. The widespread adoption of MCP by these industry leaders not only highlights its growing relevance but also points to its potential as a foundational layer in AI-powered ecosystems. As more companies integrate MCP into their operations, the protocol is set to play a central role in shaping the future of AI tool integration. Looking ahead, MCP is poised to become a key enabler of AI-driven workflows, driving more secure, scalable, and efficient AI ecosystems across industries.
Table 1. Overview of MCP ecosystem adoption.
@@
Anthropic (Claude) (Anthropic, 2024a) Full MCP support in the desktop version, enabling interaction with external tools.
AI Models and Frameworks OpenAI (OpenAI, 2025) MCP support in Agent SDK and API for seamless integration.
Baidu Maps (Maps, 2025) API integration using MCP to access geolocation services.
@@
Replit (Replit, 2025) AI-assisted development environment with MCP tool integration.
Microsoft Copilot Studio (Studio, 2025) Extends Copilot Studio with MCP-based tool integration.
Developer Tools Sourcegraph Cody (Cody, 2025) Implements MCP through OpenCTX for resource integration.
@@
Zed (Zed, 2025) Provides slash commands and tool integration based on MCP.
@@
IDEs/Editors Windsurf Editor (Editor, 2025) AI-assisted IDE with MCP tool interaction.
@@
Cloudflare (Cloudflare, 2025) Provides remote MCP server hosting and OAuth integration.
Cloud Platforms and Services Block (Square) ((2025), Square) Uses MCP to enhance data processing efficiency for financial platforms.
@@
Apify MCP Tester (Tester, 2025) Connects to any MCP server using SSE for API testing.
Web Automation and Data LibreChat (LibreChat, 2025) Extends the current tool ecosystem through MCP integration.
Goose (Goose, 2025) Allows building AI agents with integrated MCP server functionality.
@@
Anthropic has not yet released an official MCP marketplace, but the vibrant MCP community has stepped in to fill this gap by creating numerous independent server collections and platforms. As shown in Table 2, platforms such as MCP.so (mcp.so, 2025), Glama (glama.ai, 2025), and PulseMCP (et al., 2025) host thousands of servers, allowing users to discover and integrate a wide range of tools and services. These community-driven platforms have significantly accelerated the adoption of MCP by providing accessible repositories where developers can publish, manage, and share their MCP servers. Desktop-based solutions like Dockmaster (mcp dockmater, 2025) and Toolbase (gching, 2025) further enhance local MCP deployment capabilities, empowering developers to manage and experiment with servers in isolated environments. The rise of community-driven MCP server ecosystems reflects the growing enthusiasm for MCP and highlights the need for a formalized marketplace.
Table 2. Overview of MCP server collections and deployment modes (As of March 27, 2025).
@@
MCP.so mcp.so Website 4774 mcp.so
Glama glama.ai Website 3356 glama.ai
PulseMCP Antanavicius et al. Website 3164 pulsemcp.com
Smithery Henry Mao Website 2942 smithery.ai
Dockmaster mcp-dockmaster Desktop App 517 mcp-dockmaster.com
Official Collection Anthropic GitHub Repo 320 modelcontextprotocol/servers
AiMCP Hekmon Website 313 aimcp.info
MCP.run mcp.run Website 114 mcp.run
Awesome MCP Servers Stephen Akinyemi GitHub Repo 88 appcypher/mcp-servers
mcp-get registry Michael Latman Website 59 mcp-get.com
Awesome MCP Servers wong2 Website 34 mcpservers.org
OpenTools opentoolsteam Website 25 opentools.com
@@
make inference mkinf Website 20 mkinf.io
Awesome Crypto MCP Servers Luke Fan GitHub Repo 13 badkk/crypto-mcp-servers
@@ 4.2.1. OpenAI: MCP Integration in AI Agents and SDKs
OpenAI has adopted MCP to standardize AI-to-tool communication, recognizing its potential to enhance integration with external tools. Recently, OpenAI introduced MCP support in its Agent SDK, enabling developers to create AI agents that seamlessly interact with external tools. In a typical workflow, developers use the Agent SDK to define tasks that require external tool invocation. When an AI agent encounters a task like retrieving data from an API or querying a database, the SDK routes the request through an MCP server. The request is transmitted via the MCP protocol, ensuring proper formatting and real-time response delivery to the agent. OpenAI’s plan to integrate MCP into the Responses API will streamline AI-to-tool communication, allowing AI models like ChatGPT to interact with tools dynamically without extra configuration. Additionally, OpenAI aims to extend MCP support to ChatGPT desktop applications, enabling AI assistants to handle various user tasks by connecting to remote MCP servers, further bridging the gap between AI models and external systems.
@@ 4.2.2. Cursor: Enhancing Software Development with MCP-Powered Code Assistants
Cursor uses MCP to enhance software development by enabling AI-powered code assistants that automate complex tasks. With MCP, Cursor allows AI agents to interact with external APIs, access code repositories, and automate workflows directly within the integrated development environment. When a developer issues a command within the IDE, the AI agent evaluates whether external tools are needed. If so, the agent sends a request to an MCP server, which identifies the appropriate tool and processes the task, such as running API tests, modifying files, or analyzing code. The results are then returned to the agent for further action. This integration helps automate repetitive tasks, minimizing errors and enhancing overall development efficiency. By simplifying complex processes, Cursor boosts both productivity and accuracy, allowing developers to execute multi-step operations effortlessly.
@@ 4.2.3. Cloudflare: Remote MCP Server Hosting and Scalability
Cloudflare has played a pivotal role in transforming MCP from a local deployment model to a cloud-hosted architecture by introducing remote MCP server hosting. This approach eliminates the complexities associated with configuring MCP servers locally, allowing clients to connect to secure, cloud-hosted MCP servers seamlessly. The workflow begins with Cloudflare hosting MCP servers in secure cloud environments that are accessible via authenticated API calls. AI agents initiate requests to the Cloudflare MCP server using OAuth-based authentication, ensuring that only authorized entities can access the server. Once authenticated, the agent dynamically invokes external tools and APIs through the MCP server, executing tasks such as data retrieval, document processing, or API integration. This approach not only reduces the risk of misconfiguration but also ensures seamless execution of AI-powered workflows across distributed environments. Furthermore, Cloudflare’s multi-tenant architecture allows multiple users to securely access and manage their own MCP instances, ensuring isolation and preventing data leakage. Cloudflare’s solution thus extends MCP’s capabilities by enabling enterprise-grade scalability and secure multi-device interoperability.

### v3

@@ 3.3.4. Operation Phase
The operation phase represents the runtime stage of the MCP server lifecycle, where the deployed server actively interacts with users, clients, and external resources. During this phase, the server interprets user intent, mediates resource access, invokes registered tools, and manages ongoing sessions to deliver reliable and contextual responses. Intent analysis is the initial operation stage in which user inputs are parsed, contextualized, and mapped to the most suitable MCP capabilities. The host or client forwards user requests to the MCP server, where intent interpretation logic determines whether to trigger a tool, retrieve a resource, or engage a prompt. Accurate intent analysis directly affects usability and effectiveness. Misinterpretation can lead to incorrect tool execution or unnecessary resource calls, potentially increasing latency or amplifying risk exposure. External resource access occurs when the server needs to obtain supplementary data from third‑party systems or knowledge repositories. Such access is handled via predefined resource interfaces and is strictly governed by security permissions defined during creation. Each access request must comply with authentication, authorization, and sandboxing policies to prevent data leakage or external dependency compromise. Tool invocation is the execution stage where the MCP server or client triggers a registered tool according to the interpreted intent. Each tool invocation includes structured parameter passing, execution monitoring, and result serialization before returning structured outputs to the user session. Reliability and isolation at this stage determine runtime stability. Session management maintains the logical continuity between user interactions and server processes. It includes establishing, monitoring, and closing session contexts that connect the MCP host, client, and user interface.
3.3.5. Maintenance Phase
After deployment and operation, an MCP server must be periodically maintained to ensure its reliability, security, and compliance with evolving system requirements. Version control ensures that all updates to the MCP server, including bug fixes and capability extensions, are tracked through an auditable revision system. Proper version management enables rollback to stable releases, facilitates controlled feature rollout, and supports compatibility testing across client environments. Configuration change management governs modifications to runtime parameters, environmental variables, or authorization policies. All configuration adjustments should follow a controlled workflow, such as change‑request approval and pre‑deployment validation, to prevent inadvertent disruptions. Access audit records and reviews all authentication and authorization events related to the MCP server. This includes tracking user sessions, privilege escalations, and external resource permissions utilized during operation. Regular access auditing allows administrators to identify abnormal patterns, enforce least‑privilege policies, and comply with organizational or regulatory security frameworks. Log audit focuses on continuous collection and analysis of operational logs generated by the server, tools, and clients. Centralized log aggregation and integrity‑protected storage support forensic traceability and incident response. Automated log analysis, through anomaly detection or correlation with known attack signatures, can reveal early indicators of compromise or performance degradation. Effective log auditing transforms runtime data into actionable intelligence for system governance.
@@ 4.1. Ecosystem Overview
Table 1 demonstrates how MCP has gained significant traction across diverse sectors, signaling its growing importance in enabling seamless AI-to-tool interactions. The dataset summarized in this table was compiled through manual inspection, drawing on official documentation from the earliest MCP supporters and extended via community discussions and repository mining. We primarily included mature products and companies with verifiable MCP integration. Although this collection provides a representative snapshot of the current MCP landscape, we acknowledge that it is not exhaustive. To enhance transparency and facilitate future updates, the dataset will be maintained as a public repository [^1] https://github.com/security-pride/MCP_Landscape, enabling ongoing community contributions and periodic verification.
Table 1. Overview of MCP ecosystem adoption (updated to Sept. 2025).
@@
AI Models and Frameworks Anthropic (Claude) (Anthropic, 2024a) Full MCP support in the desktop version, enabling interaction with external tools.
OpenAI (OpenAI, 2025b) MCP support integrated across products and within the Agent SDK for seamless interoperability.
Google DeepMind (DeepMind, 2025) Added MCP protocol support for the Gemini model family, enabling standardized tool invocation.
Baidu Maps (Maps, 2025) API integration using MCP to access geolocation and spatial services.
@@
Developer Tools Replit (Replit, 2025) AI-assisted development environment with integrated MCP tools.
Microsoft Copilot Studio (Studio, 2025) Officially announced MCP support in March 2025, enabling unified tool integration.
Sourcegraph Cody (Cody, 2025) Implements MCP through OpenCTX for resource integration.
@@
IDEs/Editors Zed (Zed, 2025) Provides slash commands and tool integration based on MCP.
@@
Windsurf Editor (Editor, 2025) AI-assisted IDE with MCP tool interaction.
@@
Cloud Platforms and Services Cloudflare (Cloudflare, 2025) Provides remote MCP server hosting, OAuth integration, and scalable multi-tenant infrastructure.
Tencent Cloud (Cloud, 2025c) Added MCP plugin and transport support, enabling AI SDK-level integration.
Alibaba Cloud Bailian (Cloud, 2025a) Introduced full lifecycle MCP service for large-scale batch integration.
Huawei Cloud (Cloud, 2025b) Launched AI-native run platform with built-in MCP module to accelerate trusted AI deployment.
Block (Square) ((2025), Square) Uses MCP to enhance data processing efficiency for financial platforms.
@@
Alipay / Ant Group (Alipay, 2025; Group, 2025) Released “Alipay MCP Server” and “MCP Zone”, offering unified payment and tool orchestration.
Web Automation and Data Apify MCP Tester (Tester, 2025) Connects to any MCP server using SSE for API testing.
LibreChat (LibreChat, 2025) Extends the current tool ecosystem through MCP integration.
Baidu Create Conference (Baidu, 2025) Established a dedicated MCP ecosystem forum to promote developer collaboration.
Notably, leading AI companies such as Anthropic (Anthropic, 2024a) and OpenAI (OpenAI, 2025b) have integrated MCP to enhance agent capabilities and improve multi-step task execution. This adoption by industry pioneers has set a precedent, encouraging other major players to follow suit. Chinese tech giants like Baidu (Maps, 2025) have also incorporated MCP into their ecosystems, highlighting the protocol’s potential to standardize AI workflows across global markets. Developer tools and IDEs, including Replit (Replit, 2025), Microsoft Copilot Studio (Studio, 2025), JetBrains (JetBrains, 2025), and TheiaIDE (TheiaAI/TheiaIDE, 2025), leverage MCP to facilitate agentic workflows and streamline cross-platform operations. This trend indicates a shift toward embedding MCP in developer environments to enhance productivity and reduce manual integration efforts. Furthermore, cloud platforms like Cloudflare (Cloudflare, 2025) and financial service providers such as Block (Square) ((2025), Square) and Stripe (Stripe, 2025) are exploring MCP to improve security, scalability, and governance in multi-tenant environments. The widespread adoption of MCP by these industry leaders not only highlights its growing relevance but also points to its potential as a foundational layer in AI-powered ecosystems. As more companies integrate MCP into their operations, the protocol is set to play a central role in shaping the future of AI tool integration. Looking ahead, MCP is poised to become a key enabler of AI-driven workflows, driving more secure, scalable, and efficient AI ecosystems across industries.
While these examples show clear momentum, the overall MCP ecosystem remains uneven in maturity. Many community-hosted servers are small experiments or early prototypes. Only a limited number demonstrate strong reliability, stable maintenance, or standardized documentation. A deeper issue lies in platform governance. Many large internet platforms operate as walled gardens. They treat data as a controlled resource and limit access to interfaces or user behavior data to maintain a competitive advantage. In such systems, the open and composable design that MCP promotes may face natural resistance. We note that platforms might only expose low-risk or secondary functions through MCP connections. For example, an online map service might allow MCP access to public location data or route information, but not to personalized navigation history or user-generated content. Similarly, a social media platform could open an MCP interface for basic profile lookup or public post retrieval, while keeping private messages and recommendation algorithms restricted. These partial integrations show that platform openness often depends on strategic and privacy considerations rather than technical feasibility. This reflects a tension between open interoperability and proprietary data control. The success of MCP will partly depend on whether large platforms are willing to share meaningful data and support cross-system collaboration.
@@
Although Anthropic has not yet released an official MCP marketplace, the broader community has actively filled this gap by creating numerous independent MCP server collections and discovery platforms. As summarized in Table 2, we identified and catalogued all publicly available MCP server directories accessible as of September 2025. These include a diverse range of deployment mode, including websites, GitHub repositories, and desktop applications, demonstrating how third-party developers are collectively working to establish a complete and sustainable MCP ecosystem. The data presented in Table 2 were collected and verified manually. We began by surveying known MCP-related repositories on GitHub and popular community forums, followed by direct examination of listed directory or application to confirm its activity, accessibility, and declared number of hosted servers. Most platforms explicitly report total server counts, which we cross-checked against their public listings; where such data were missing, we manually counts from available catalog pages. This collection process yielded a consolidated dataset encompassing 26 major MCP collections. Platforms such as MCP.so (mcp.so, 2025), Glama (glama.ai, 2025), and PulseMCP (et al., 2025) host thousands of servers, allowing users to discover and integrate a wide range of tools and services. These community-driven platforms have significantly accelerated the adoption of MCP by providing accessible repositories where developers can publish, manage, and share their MCP servers. Desktop-based solutions like Dockmaster (mcp dockmater, 2025) and Toolbase (gching, 2025) further enhance local MCP deployment capabilities, empowering developers to manage and experiment with servers in isolated environments.
Despite this rapid growth, the quality of community-maintained MCP markets remains uneven. Platforms like MCP.so host thousands of entries but lack formal security or identity verification mechanisms. To assess reliability, we randomly sampled 300 servers listed on MCP.so. Among them, 30 contained the term “MCP” in the project title but did not refer to the Model Context Protocol, and 18 were in active development or unavailable at the time of inspection. These findings suggest that large community directories may overstate their effective numbers, and server quality varies widely across listings. In contrast, platforms such as mcp-get implement verification and signing processes, but the number of verified servers remains very small and user adoption is limited. The overall landscape therefore reflects a trade‑off between openness and quality control: while community initiatives accelerate ecosystem growth, they also highlight the need for stronger validation standards and sustainable curation practices.
Table 2. Overview of MCP server collections and deployment modes (As of Sept. 2025).
@@
MCPWorld Baidu Website 26,404 mcpworld.com
MCP.so mcpso Website 16,592 mcp.so
MCP Servers Repository mcprepository Website 13,596 mcprepository.com
AIbase MCP AIbase Website 12,448 mcp.aibase.com
Glama glama.ai Website 9,415 glama.ai
Smithery Henry Mao Website 6,888 smithery.ai
PulseMCP Antanavicius et al. Website 6,072 pulsemcp.com
ModelScope MCP Marketplace modelscope Website 5,441 modelscope.cn/mcp
Awesome MCP Servers wong2 Website 2,402 mcpservers.org
Cursor Directory MCP Cursor Website 1,800 cursor.directory/mcp
Official Collection1 Anthropic GitHub Repo 1,204 modelcontextprotocol/servers
AiMCP Hekmon Website 907 aimcp.info
Dockmaster mcp-dockmaster Desktop App 516 mcp-dockmaster.com
MCP Market mcpmarket.com Website 463 mcpmarket.com
MCP.run mcp.run Website 242 mcp.run
Awesome MCP Servers Stephen Akinyemi GitHub Repo 217 appcypher/mcp-servers
CLine MCP Marketplace Cline Website 154 cline.bot/mcp-marketplace
Bailian MCP Market Aliyun Website 151 bailian.console.aliyun.com
OpenTools opentoolsteam Website 148 opentools.com
Awesome Remote MCP Servers JAW9C GitHub Repo 79 jaw9c/awesome-remote-mcp-servers
MCP Server Hub mcpserverhub Website 71 mcpserverhub.com
mcp-get Michael Latman Website 59 mcp-get.com
MCP Marketplace Higress.ai Website 50 mcp.higress.ai
@@
make inference mkinf Website 23 mkinf.io
Awesome Crypto MCP Servers Luke Fan GitHub Repo 12 badkk/crypto-mcp-servers
  • 1
    Official Collection refers to the list of MCP servers curated by Anthropic. Anthropic has also launched the community-driven MCP Registry, a preview service that provides a centralized directory for discovering MCP servers (https://github.com/modelcontextprotocol/registry).
@@ 4.2.1. OpenAI: MCP Integration in AI Agents and SDKs
We chose OpenAI as the case study because it holds significant influence in the field of agent development. The Agents SDK (OpenAI, 2025a), as a mainstream development framework, covers a wide range of real-world applications. It incorporates basic capabilities such as managing conversations, invoking tools, task delegation, input and output verification, and session recording. Developers only need to write a small amount of Python code to enable the intelligent agent to handle multiple rounds of conversations automatically, flexibly call various tools, and even allow multiple intelligent agents to collaborate to complete complex tasks. After the MCP was added, the Agents SDK became more efficient in terms of tool integration. In the past, developers had to customize the connection logic separately for each new tool or external service. This was not only time-consuming but also prone to errors. Now, with the help of MCP, developers only need to configure the address of the MCP tool, and the agent can automatically discover, connect to, and call it, regardless of where these tools are deployed. The entire integration process has become more unified, and maintaining and extending new tools has become much simpler. ChatGPT currently supports MCP in developer mode. Users can directly connect MCP tools within ChatGPT, enabling them to do more than just query data. They can also write to external systems, trigger automated tasks, or link multiple tools together to perform complex operations. The introduction of MCP has transformed ChatGPT from a past question-and-answer assistant into a platform that can continuously expand its capabilities.
During the process of promoting and practicing MCP, OpenAI has accumulated rich experience and provided many applicable practices for the industry. Through this case, we can observe how MCP helps developers avoid detours and accelerate the transformation of ideas into practical applications. This not only makes the integration of tools smoother but also makes AI systems more open and flexible.
@@ 4.2.2. Cursor: Enhancing Software Development with MCP-Powered Code Assistants
In recent years, AI-assisted programming tools have significantly simplified the software development process. Tools like Cursor, Claude Code, and Cline enable developers to interact with the system using natural language. These assistants can automatically generate code, assist with debugging, and also perform structural reconfiguration. As a result, developers can focus more on business logic, leading to a significant improvement in development efficiency and code quality. The integration of the MCP protocol has further expanded the capabilities of AI-assisted programming tools. Taking Cursor as an example, MCP enables it to directly access external APIs, code repositories, and various automation tools. In practical applications, this process typically includes steps such as instruction parsing, tool scheduling, task distribution, and result feedback. Developers simply input tasks in Cursor, such as “Help me perform a page visit and take a screenshot on qq.com”, and Cursor will analyze whether external tool support is needed and request services from the backend through the MCP protocol. The MCP server is responsible for scheduling the appropriate tools, such as Playwright MCP, to automatically complete tasks like page visits and screenshots, and return the results to the agent, ultimately presenting them to the developer. Even for more complex automated testing tasks, such as form submissions, AJAX requests, or multi-page navigation, developers do not need to leave the IDE environment. All processes can be automated and standardized.
The introduction of MCP has greatly enhanced the flexibility and scalability of Cursor. Developers no longer need to write repetitive adaptation codes for each new tool; they only need to declare the tool addresses to access different types of services, regardless of whether these tools are running locally, remotely, or in the cloud. This mechanism reduces the difficulty of expansion and maintenance. Cursor can also automatically coordinate different tools based on the current development scenario, helping developers complete multiple operations. As a result, the development experience becomes smoother and the efficiency is significantly improved. Of course, the integration of MCP also brings some challenges. For example, the data compatibility between different tools needs to be addressed, and permission management and security isolation are even more important. Distributed calls may cause network delays, which have an impact on the overall experience. Despite these issues, after integrating MCP, the intelligence and openness of the development environment have significantly improved. This case helps us understand the profound impact of MCP on actual software development and provides a reference for the future development of the industry.
@@ 4.2.3. Cloudflare: Remote MCP Server Hosting and Scalability
Cloudflare transforms MCP from a local-only technology into a scalable, cloud-based solution, as illustrated in Figure 4. By hosting MCP servers in the cloud, Cloudflare removes the need for users to configure and maintain servers on their own machines. This shift lowers the technical barrier for both developers and end users. Cloudflare integrates managed authentication with OAuth 2.0, which ensures that only authorized agents and users can access MCP servers and the tools they provide. The platform also supports persistent state and secure data storage through technologies like Durable Objects and Workers KV. Each MCP session can reliably maintain its own data, even as usage grows or as users switch devices. With these capabilities, MCP servers running on Cloudflare can connect to external APIs, automate multi-step workflows, and interact with a broad range of third-party services. Developers can focus on building useful features, while Cloudflare handles critical concerns such as security, scaling, and network connectivity. This multi-tenant design allows different users or organizations to safely share the same infrastructure without the risk of data leakage.
Figure 4. Architecture of local and remote MCP communication via proxy.
Deploying MCP servers on a cloud platform introduces certain challenges. Developers must carefully define authorization scopes to limit access to sensitive resources. They also need to address privacy and data residency requirements for users in different regions. Real-time agent workflows may require special attention to maintain low latency and stable connections. Cloudflare provides tools to address many of these concerns, but secure tool design and session management remain important responsibilities for developers. Moving MCP to the cloud with Cloudflare extends the value of the protocol to a much wider audience. Both technical and non-technical users can now access secure, scalable AI-powered tools from any device. This cloud-native approach helps MCP become a universal interface for safe, extensible, and cross-device AI automation. Cloudflare’s platform shows how modern cloud infrastructure can make advanced AI agent ecosystems accessible and practical in real-world applications.